### v1

@@ 3 Physical Picture
$H=\mu\sum_{i}n_{i\sigma}-t\sum_{i\neq j\sigma}^{n.n}c^{\dagger}_{i\sigma}c_{j\sigma}+\frac{U}{2}\sum_{i\sigma}n_{i\sigma}n_{i-\sigma},$
@@ 6.1 Free Electron Gas- The Perfect Quantum gas
To repeat, this Hamiltonian corresponds to a non-interacting system (U=0) as compared to the Hamiltonian described in the physical picture in Sec.(3). Noting that for fermions, $\eta=-1$, we define Green’s function for this case as follows.
@@ 6.2 Phonon Gas
7 Conclusion
We have solved the Hubbard Hamiltonian in atomic limit, an exact solution and some approximate methods such as Hartree Fock[2], Hubbard I[2] and Hubbard III[3]. These articles employ a semicircular density of states. We have extended it to include delta function, toy semiconductor model and rectangular density function. This work is included in the lecture notes on ”Model Hamiltonians solved by Zubarev’s Double Time Green’s Function” .

### v2

@@ Abstract
These lecture notes present a comprehensive and powerful many-body technique pioneered in 1960 by D. N. Zubarev. The technique, known as the Zubarev Double Time Green’s Function method, was used extensively by leading solid state physicists such as John Hubbard and Laura Roth in the 1960s. We present the technique and apply it to the non-interacting electron and boson gas. We next consider the (many-body) Hubbard model and show how it yields the Stoner criterion for ferromagnetism. It is easily extendable to superconductivity and related problems. Our treatment is pedagogical and understandable to those with just an elementary understanding of second quantization.
@@ 1.2 Physical Picture
$H=\mu\sum_{i\sigma}n_{i\sigma}-t\sum_{i\neq j\sigma}^{n.n}c^{\dagger}_{i\sigma}c_{j\sigma}+\frac{U}{2}\sum_{i\sigma}n_{i\sigma}n_{i-\sigma},$
@@ 4.1 Free Electron Gas- The Perfect Quantum gas
To repeat, this Hamiltonian corresponds to a non-interacting system (U=0) as compared to the Hamiltonian described in the physical picture in Sec.(1.2). Noting that for fermions, $\eta=-1$, we define Green’s function for this case as follows.
@@ 4.2 Phonon Gas
5 Acknowledgment
SS prepared these notes under the guidance of VS as part of a semester project in 2018 at UM-DAE CEBS. VS acknowledges support from the Raja Ramanna Fellowship by the DAE. SS acknowledges support from the INSPIRE scholarship for funding her undergraduate studies at UM-DAE CEBS.